### v1

@@ Abstract
Generation of squeezed light is usually implemented in nonlinear or materials. Semiconductor lasers and optical amplifiers (SOAs) also offer nonlinearities but they differ from passive elements in that they add amplified spontaneous emission noise (ASE). In a semiconductor laser, squeezing to below the shot noise limit has been demonstrated. An SOA contains no cavity and it adds significant noise. Gain saturation can lead, in principle, to squeezing of the photon number quadrature to below the shot noise level but often the noise is reduced only to below the ASE level of a linear amplifier. At the same time, the noise in the phase quadrature increases according to the Heisenberg uncertainty principle. Short resonant pulses interacting with a quantum dot SOA induce coherent effects such as Rabi oscillations. Here, we demonstrate, for the first time, that Rabi oscillations cause cyclical noise squeezing which varies periodically with the excitation pulse area. The noise in the present experiments does not reach the quantum limit so we term this condition quasi-squeezing. It occurs during the portions of the Rabi cycle when the quantum dots provide gain and repeats with every fourfold increase of the pulse excitation energy which amounts to a 2 increase in pulse area. In all other cases, the noise exhibits the properties of a coherent state.
@@
squeezing, quantum dot, Rabi oscillations, semiconductor optical amplifier1 Introduction
@@
Squeezed light is a non classical state of an optical signal in which the uncertainty (noise) in one quadrature (for example the photon number X) describing the field is reduced at the expense of the other quadrature (the phase P), while obeying the Heisenberg uncertainty [1, 2, 3]. Squeezed optical signals allow for improved fiber communication [4, 5, 6], and are imperative for precise measurements in quantum sensing and metrology [7], as well as for continuous variable quantum communication [8, 9]. Generation of squeezed light requires a strong nonlinearity which is most often implemented in passive or materials [10, 11, 12, 13, 14].
@@
Strong nonlinearities are also available in optically active media, such as lasers and semiconductor optical amplifiers (SOAs). In the context of light squeezing, active media differ from passive materials in that they add spontaneous emission noise. The nonlinearity in an active device, and its consequent ability to squeeze the noise, stems from saturation which is determined, in turn, by the dynamical properties. In solid state and doped fiber devices, dynamics are slow and saturation is governed by average powers. In contrast, the dynamics of semiconductor devices are fast, and this has a major impact on the noise properties under saturated conditions [15, 16, 17].
@@
Squeezing in a DC driven semiconductor laser to below the quantum limit was demonstrated by Y. Yamamoto’s group several decades ago [18, 19]. Similar squeezing levels were achieved in A. Yariv’s group by applying a weak dispersive feedback to the laser [20]. A model for squeezing in semiconductor lasers was reported in [21]. The noise of a saturated SOA was shown both theoretically and experimentally to exhibit a spectral hole in the vicinity of the saturating signal [22, 15, 23]. Clear narrowing of the the statistical distribution of the power spectral density in a saturated SOA was demonstrated experimentally for continuous waves (CW) [24] as well as pulse [25] excitations, and confirmed by a simulation [26].
@@
Short pulse propagation in an SOA induces several nonlinear coherent interactions [27]. Once the pulses are sufficiently intense and their duration is short in comparison to the semiconductor coherence time (which is roughly 1 ps at room temperature [28]), they induce coherent interactions such as Rabi oscillations, which were demonstrated experimentally mainly in quantum-dot (QD) SOAs [27, 29, 30, 31], but also in a quantum cascade laser [32], and were modeled for quantum well SOAs [33]. This means that during the pulse, the SOA oscillates between gain and absorption. This leaves, in turn, a direct imprint on the temporal profiles of the amplitude and phase of the SOA output pulse. The carrier density in the SOA also oscillates and this affects the amplified spontaneous emission noise accompanying the pulse.
We demonstrate here, for the first time, a direct interplay between Rabi oscillations and noise squeezing in a QD SOA operating at 1550 nm and at room temperature. In the present experiments, gain saturation does not squeeze the noise below the shot noise limit but rather reduces it, in the photon number quadrature, to below its value under linear gain conditions. We term this condition quasi-squeezing. At the same time, the noise in the other quadrature is correspondingly increased.
To induce Rabi oscillations, we excite the QD SOA with 100 fs pulses and characterize the noise of the resulting output pulses using homodyne detection, where a replica of the excitation pulse acts as the local oscillator (LO). Noise squeezing is found to occur only under conditions of gain saturation and it is cyclical with the input pulse excitation energy. Namely, squeezed noise is observed when the excitation pulse energies increase by factors of four which amounts to a 2 increase in pulse area. For other pulse excitation energies, the QD SOA output exhibits a coherent state. This periodic behavior extends over a very wide range of excitation pulse energies.
@@ 2 Results and Discussion
The experimental setup for homodyne characterization of ultra short pulses following coherent interaction with a QD SOA is shown in Fig. 1. A 100 femtosecond laser pulse is split with one part serving as the LO and the second (signal) is coupled to a QD optical amplifier. The signal pulse interacts with the QD gain medium, and the transmitted field is recombined with the LO, whose delay is controlled via an optical delay line. The resulting interference is detected by a balanced photodetector, and the difference signal is recorded using an oscilloscope. Scanning the LO phase over 2 radians for about one thousand times yields a histogram of the current uncertainty. The histogram is transformed into a Wigner function using the Inverse Radon Transformation [34]. The insets in Fig. 1 show respectively the Wigner functions of the amplified spontaneous emission noise under linear conditions namely, with no input signal and of a typicsl quasi squeezed state.
@@
The homodyne system serves also as a cross-correlator where the LO pulse samples the output pulses of the QD gain medium while the relative temporal delay between them is scanned. Such cross-correlations for three different excitation pulse energies are shown in the left column of Fig. 2. The cross-correlations reveal that the 100-fs input pulse broadens to about 1 ps and attains an amplitude profile resembling the results measured using cross-correlation frequency resolved optical gating (X-FROG) [29]. Homodyne measurements were performed with the LO sampling the signal at different temporal locations. The center column of Fig. 2 describes histograms of the normalized balanced homodyne detector output current as a function of the LO phase, measured at the trailing peak (dashed line at 0 ps). The corresponding extracted Wigner function are depicted in the right column, revealing a clear dependence on the input pulse energy. At 6.3 and 25 pJ, the Wigner function is that of a coherent state. However, at 12.5 pJ it exhibits clear quasi-squeezing. At any other temporal position, the Wigner functions for all input pulse energies exhibit coherent states namely, they are not affected by by the QD dynamics stemming from the Rabi oscillations.
The dependence on pulse excitation energy was examined in more detail as described in Fig. 3. The upper left panel shows, for reference, the Wigner function of the amplified spontaneous emission (ASE) noise obtained with no input signal. Quasi-squeezing is found to be cyclical with a periodicity of a fourfold increase in energy, from 3.96 pJ to 15.8 pJ corresponding to a pulse area change of 2. The two quasi squeezed Wigner functions at 3.96 pJ and 15.8 pJ exhibit a photon number uncertainty which is smaller than that of the ASE as seen by the dashed circles. Note that the absolute values of the pulse area cannot be determined precisely, since a QD SOA is an inhomogeneously broadened, distributed gain medium. The pulse area is defined as where is the dipole matrix element and is the envelope of the electric field at position . According to the McCall–Hahn area theorem [35], which was originally derived for resonant two-level systems under the assumptions of negligible dephasing, no inhomogeneous broadening, and no nonlinear gain dynamics, the pulse area in an amplifying medium evolves as where is the gain coefficient. In a QD amplifier, these assumptions are only qualitatively satisfied: the ensemble exhibits inhomogeneous broadening, nonlinearity due to gain saturation, and complex carrier dynamics beyond the two-level approximation. While the exact initial pulse area cannot be determined accurately, changes in pulse area of are nevertheless exact. Quasi-squeezing occurs only when the system is in the gain regime during the Rabi cycle, while in every other case, the amplifier output is a coherent state whose Wigner function size is similar to that of the amplified spontaneous emission noise.
The cyclical nature of the quasi squeezing is supported by a simple phenomenological calculations which analyzes the spontaneous emission of an ideal two level system driven by a short excitation pulse , whose area is defined as . The pulse induces Rabi oscillations at a frequency , where is the transition dipole moment and is the reduced Planck constant. In the rotating frame of the excitation pulse, the instantaneous excited-state population evolves as:
Spontaneous emission noise is incorporated by adding Gaussian noise whose variance is proportional to the excited-state population. The output electric field becomes:
Both noise terms, and take on the form of a normal distribution
where is the spontaneous emission rate and the two noise terms are related to each other via the Henry’s linewidth enhancement parameter [36].
To estimate the resulting quantum state, we compute the average field amplitude and phase from the noisy output: . This complex amplitude is scaled and used to construct a coherent density matrix . The Wigner function is then calculated directly from this density matrix using the standard QuTiP routine: , where denotes the Wigner function of state .
Simulated Wigner functions are shown in Fig. 4 for input pulse areas ranging from to 5. The periodical nature of quasi-squeezing is clearly observed for pulse areas that differ by 2, consistent with the measured data.
3 Conclussion
@@
In this work, we demonstrated that a resonant ultra-short pulse excitation of a QD SOA induces cyclical quasi-squeezing, driven by coherent Rabi oscillations. Specifically, we observed that the output photon number quadrature noise is periodically reduced below the ASE level of a linear amplifier, while the conjugate phase quadrature noise increases correspondingly. This quasi-squeezing effect repeats with every fourfold increase in the excitation pulse energy, equivalent to a 2 increment in pulse area.
The underlying mechanism arises from the periodic evolution between gain and absorption in the QD SOA during Rabi oscillations. The noise is squeezed only during portions of the Rabi cycle when the QDs provide gain; otherwise, the output exhibits the noise properties of a coherent state. This behavior is in qualitative agreement with predictions based on the McCall–Hahn area theorem, despite the inhomogeneous broadening and complex carrier dynamics in QD SOAs.
@@
While in the present experiments the noise does not reduce to below the quantum shot noise limit, it represents a clear manipulation of amplifier-added noise via coherent light–matter interactions. This demonstrates a fundamentally new regime of amplifier operation that could enable engineered noise profiles in active photonic devices. Potential applications include tailored noise suppression in optical amplifiers for quantum communication systems, or enhanced sensitivity in measurement schemes where reduced amplitude noise, is crucial.
@@
Future research should explore extending this effect to other material systems, for example perovskite based amplifiers, and investigate whether cavity enhancement or tailored feedback could ensure squeezing below the shot noise limit. Additionally, incorporating pulse shaping techniques to precisely control the temporal dynamics of the Rabi oscillations [37] may open new avenues for deterministic quantum noise engineering on chip-scale photonic platforms.
@@ 4.1 Experimental setup
The experimental setup consists of a femtosecond fiber-based laser (Toptica FemtoFiber pro) operating at a center wavelength of 1550 nm, delivering pulses with a duration of approximately 100 fs. The optical path length was controlled using a motorized optical delay line (Physik Instrumente, model V-508, step resolution of 2 nm) to achieve precise temporal scanning. The output beams were combined and directed to a balanced photodetector (Femto, model HBPR-100M-60K-IN-FST) for detection of the interference signal. The resulting photocurrent was recorded using an oscilloscope (Keysight, model EXR404A, with 4 GHz bandwidth) for subsequent analysis. The delay line was moving periodically at rate of 5 Hz and an amplitude of 2 m and correspondingly produced a position signal that was sampled by the oscilloscope in sync with the signal from the balanced detector.
@@ 4.2 QD SOA fabrication
The active region of the gain medium we used comprised six layers of high-density cm-2 InAs QDs grown by molecular beam epitaxy in the Stranski - Krastanow mode. The optical amplifier is fabricated as a 2 wide, 1.5 mm long ridge waveguide whose end facets were anti-reflection coated. The QDs exhibit a narrow photo-luminescence linewidth, measured at 10 K, of 17 meV for a single layer and 26 meV for the six-layer stack. This linewidth represents a record uniformity compared to all reported QDs in any material system. Details of the device fabrication and characterization are given in [38]. The emission power of the measured electro-luminescence spectra increases with applied bias but the spectral shape remains unchanged [39].
@@ 4.3 Wigner function analysis
Histograms of the current at the balanced detector output, measured with a real time oscilloscope, is constructed from one thousand, periodic scans of the delay line with an amplitude of 2m. Wigner functions are obtained from those histograms using the inverse Radon transform. This transform is intended for continuous waves (CW) or narrow-spectrum signal. The spectral width of the 100 fs excitation pulses is roughly 3% of the optical frequency and hence, the assumption of narrow band signal is valid.
Acknowledgements The reearch was partially supported by the Israel Science Foundation under grant number 460/21 and by the Hellen Diller Quantum Center at Technion. IK acknowledges financial support from the Ministry of Innovation, Science and Technology of Israel.

### v2

@@ Abstract
Squeezed light is usually generated using passive nonlinear materials. Semiconductor lasers and optical amplifiers (SOAs) also offer nonlinearities but they differ in that they add amplified spontaneous emission (ASE). Squeezing to below the vacuum level has been demonstrated in a semiconductor laser, and gain saturation in SOAs can likewise reduce photon-number fluctuations to, and in some cases below, the vacuum limit. Here, we demonstrate that Rabi oscillations in room-temperature quantum-dot SOAs, induced by short resonant pulses, cause cyclical noise modification that repeat with every change of 2 in pulse area, corresponding to a fourfold increase in excitation pulse energy. Homodyne measurements reveal in those cases elliptical Wigner functions corresponding to squeezed thermal states and in certain regimes, the state is squeezed to below the vacuum level. At other pulse areas, the Wigner functions are circular representing thermal coherent states. This periodic behavior persists over two orders of magnitude in input pulse energy, spanning several 2 cycles. Under specific bias and excitation conditions, we further observe a non-Gaussian Wigner function featuring two bright lobes. Although its precise nature remains unresolved, this structure may be consistent with a Schrödinger cat–like state whose accompanying negativity is suppressed due to an approximately 10 dB optical output loss. Notably, the emergence of this non-Gaussian state is itself periodic in excitation pulse energy.
@@
noise squeezing, quantum dots, Rabi oscillations, semiconductor optical amplifier1 Introduction
@@
Squeezed light is a non-classical state of an optical signal in which the uncertainty in one quadrature (for example, the photon number) is reduced at the expense of the other quadrature (the phase), while obeying the Heisenberg uncertainty [1, 2, 3]. Squeezed optical signals allow for improved fiber communication [4, 5, 6], and are imperative for precise measurements in quantum sensing and metrology [7], as well as for continuous variable quantum communication [8, 9]. Generation of squeezed light requires a strong nonlinearity which is most often implemented in passive or materials [10, 11, 12, 13, 14].
@@
Strong nonlinearities are also available in optically active media, such as semiconductor lasers and optical amplifiers (SOAs). Active media differ from passive materials in that they add amplified spontaneous emission noise (ASE). The nonlinearity in an active device, and its consequent ability to squeeze the noise, stems from saturation which is determined, in turn, by the medium dynamical properties. The dynamics of semiconductor devices are fast, and this has a major impact on the noise properties under saturated conditions [15, 16, 17].
@@
Squeezing in a DC driven semiconductor laser was demonstrated by Y. Yamamoto’s group several decades ago [18, 19]. Similar squeezing levels were achieved in A. Yariv’s group by applying a weak dispersive feedback to the laser [20]. A recent paper [41] reported squeezing in a quantum dot (QD) laser driven by an ultra-low-noise current source. A model for squeezing in semiconductor lasers was described in [21]. The noise of a saturated SOA was shown both theoretically and experimentally to exhibit a spectral hole in the vicinity of the saturating signal [22, 15, 23]. Clear narrowing of the the statistical distribution of the power spectral density in a saturated SOA was demonstrated experimentally for continuous waves (CW) [24] as well as pulse [25] excitations, and confirmed by a simulation [26].
@@
Ultra short pulse propagation in an SOA induces several nonlinear coherent interactions [27]. Once the pulses are sufficiently intense and their duration is short in comparison to the semiconductor coherence time (which is roughly 1 ps at room temperature [28]), they induce coherent interactions such as Rabi oscillations, which were demonstrated experimentally mainly in QD SOAs [27, 29, 30, 31], but also in a quantum cascade laser [32], and were modeled for quantum well SOAs [33]. This means that during the pulse, the SOA oscillates between gain and absorption, and this leaves a direct imprint on the temporal profiles of the amplitude and phase of the SOA output pulse. The carrier density in the SOA also oscillates and this affects the amplified spontaneous emission noise accompanying the pulse.
We demonstrate here, for the first time, a direct interplay between Rabi oscillations and noise in a QD SOA operating at 1550 nm and at room temperature. To induce Rabi oscillations, we excite the QD SOA with resonant 100 fs pulses and characterize the noise of the resulting output pulses using homodyne detection, where a replica of the excitation pulse acts as the local oscillator (LO). During the portions of the Rabi cycle when the quantum dots provide gain, the uncertainty in the photon number is reduced yielding an elliptical Wigner function. We highlight three consequent results. First, the noise modification is cyclical with the pulse excitation energy repeating with every fourfold increase of the pulse energy which amounts to a 2 increase in pulse area namely, a full cycle on the Bloch sphere. Some of these elliptical Wigner functions represent squeezed states and others describe squeezed coherent thermal states [34]. Other pulses areas yield circular Wigner functions describing coherent thermal states [35] resulting from the addition of the SOA ASE. This periodic behavior extends over a very wide range of excitation pulse energies, from 1.58 pJ to 158 pJ. Second, under specific conditions, the photon number uncertainty is clearly squeezed to below the vacuum level. Finally, in some cases, the obtained Wigner function is non-Gaussian comprising two separated bright lobes. The nature of this Wigner function is unclear as it does not represent any known quantum state. It may represent however, a Schrödinger cat-like state where the corresponding negativity is not observable due to a 10 dB output optical loss. The observation of this non-Gaussian state is also cyclical, recurring with a 2 increase in pulse area.
@@ 2 Results and Discussion
The experimental setup for homodyne characterization of ultra short pulses following coherent interaction with a room-temperature QD SOA is shown in Fig. 1. A 100 fs laser pulse is split with one part serving as the LO and the second is coupled to a QD optical amplifier. Following the interaction with the QD SOA, the transmitted field (signal) is recombined with the LO, whose relative phase is controlled via an optical delay line. The resulting interference is detected by a balanced photodetector, and the difference signal is recorded using an oscilloscope. Scanning the LO phase over 2 radians (a periodic motion over ) for about one thousand times yields a histogram of the current uncertainty. The histogram is transformed into a Wigner function using the inverse Radon transform [36]. The insets in Fig. 1 show respectively, the Wigner functions of the ASE noise under linear conditions namely, with no input signal and of a typical squeezed state.
@@
The homodyne system serves also as a cross-correlator where the LO pulse samples the signal pulses after propagating through the QD gain medium while the relative temporal delay between them is scanned. Such cross-correlations for three different excitation pulse energies are shown in the left column of Fig. 2. The cross-correlations reveal that the 100 fs excitation pulse broadens to about 1 ps and attains an amplitude profile resembling the results measured using cross-correlation frequency resolved optical gating (X-FROG) [29]. Homodyne measurements were performed with the LO sampling the signal at different temporal locations. The center column of Fig. 2 describes histograms of the normalized output current of the balanced photodetector as a function of the LO phase, measured at the trailing peak (dashed line at 0 ps). The corresponding extracted Wigner function are depicted in the right column, revealing a clear dependence on the input pulse energy. At 6.3 and 25 pJ, the circular Wigner functions represent thermal coherent states. However, at 12.5 pJ it exhibits a clear elliptical function characteristic of the reduced photon number uncertainty. This is a squeezed coherent thermal state. At any other temporal position, the Wigner functions for all excitation pulse energies exhibit coherent states, namely, they are not affected by the QD dynamics stemming from the Rabi oscillations.
The dependence on pulse excitation energy was examined in detail with the results shown in Fig. 3. The upper left panel shows for reference, the Wigner distributions of the vacuum and the ASE. Three Wigner functions, measured at 1.56 pJ, 6.28 pJ and 9.95 pJ, are round describing coherent thermal states. Their distribution is equal or is slightly wider than the standard deviation of the ASE which is shown in a green circle. The Wigner functions for 3.96 pJ and for 15.8 pJ (which is four ties larger) are elliptical. for the case of 3.96 pJ is equal to the standard deviation of the ASE, namely this is a squeezed thermal state. Increasing the input energy by a factor of four to 15.8 pJ yields a Wigner function whose uncertainty in the photon number is slightly, but nevertheless clearly, smaller than the standard deviation of the vacuum (shown in a red circle). It represents squeezing.
The experiments were repeated at a range of larger input energies, from 9.4 pJ to 158 pJ, under conditions of somewhat lower input coupling efficiencies to the QD SOA. The results are shown in Fig. 4.
The cyclical nature, with the same fourfold periodicity of the excitation energies is maintained for 9.4 pJ, 35.3 pJ and 141 pJ, yielding elliptical Wigner functions. in those is similar to the standard deviation of the ASE distribution (shown in green circles). Therefore, they do not represent squeezing but rather squeezed coherent thermal states. The higher energies yield multiple Rabi cycles, each contributing a different degree of saturation and hence, the cumulative effect of the noise modification is less efficient. The Wigner functions at 28.1 pJ, 56 pJ and 158 pJ are circular and describe once more coherent thermal states. Some elliptical Wigner functions in Fig. 3 and Fig. 4, in particular for the higher excitation energies, are tilted. This means that in addition to the obvious relationship between photon number and phase uncertainties, there exists a nonlinear coupling between them. This is a consequence of the power dependence of the Henry parameter [39, 38].
The QD SOA is a distributed, inhomogeneously broadened medium where an accurate value for the pulse area is hard to define even though the differences in area between successive measurements are essentially accurate. It turns out that there are some specific excitation energies that yield very unique results. The first is shown in Fig. 5. For an excitation energy of 25 pJ, the obtained Wigner function is unequivocally squeezed below the vacuum, as seen in Fig. 5 (a). Fig. 5 (b) shows a contour plot of the Wigner (in blue) and vacuum (in red) distributions. The state is clearly squeezed. Finally, under excitation energies of 44.5 pJ and 187 pJ, the obtained Wigner function is non-Gaussian as seen in Fig. 6. This Wigner function that comprises two bright lobes does not represent any known quantum state. It does resemble though a Schrödinger cat-like state except that their corresponding negativity is not observed. The negativity is possibly missing due to a large, close to 10 dB output optical loss [37] .
3 Conclusion
@@
In this work, we demonstrated a direct interplay between Rabi oscillations in a room-temperature QD SOA, induced by an ultra-short excitation pulse and noise. Specifically, Rabi oscillations cause cyclical modifications of the output photon number uncertainty. This effect repeats with every fourfold increase in the excitation pulse energy, equivalent to a 2 increment in pulse area. The noise in those cases takes on a value which is similar or slightly smaller than the ASE level of a linear amplifier. In some cases, it is actually squeezed to below the vacuum level. This behavior was observed for a two orders of magnitude range of excitation energies. Naturally, the conjugate phase quadrature noise increases correspondingly.
The underlying mechanism arises from the periodic evolution between gain and absorption in the QD SOA during Rabi oscillations. The photon number uncertainty is reduced during portions of the Rabi cycle when the QDs provide gain saturation; otherwise, the output exhibits the noise properties of a coherent thermal state. However, there are some specific excitation levels that yield very unique results. A clear squeezed state and a non-Gaussian Wigner functions were demonstrated. The latter resembles a Schrödinger cat-like state where the negativity is omitted, possibly due to a large optical loss between the SOA output facet and the homodyne system.
@@
Our findings represent a clear manipulation of noise (amplifier added ASE and noise carried by the input signal) via coherent light–matter interactions. This is a fundamentally new regime of amplifier operation that could enable engineered noise profiles in active photonic devices. Potential applications include tailored noise suppression in optical amplifiers for quantum communication systems, or enhanced sensitivity in measurement schemes where reduced amplitude noise, is crucial.
@@
Future research should explore this effect in cavity-based system with tailored feedback. Additionally, incorporating pulse shaping techniques to precisely control and optimize the temporal dynamics of the Rabi oscillations [karni2016coherent] may enhance the noise reduction while opening new avenues for deterministic quantum noise engineering on chip-scale photonic platforms.
@@ 4.1 Experimental setup
The experimental setup consists of a femtosecond fiber-based laser (Toptica FemtoFiber pro) operating at a center wavelength of 1550 nm, delivering pulses with a duration of approximately 100 fs. The optical path length was controlled using a motorized optical delay line (Physik Instrumente, model V-508, step resolution of 2 nm) to achieve precise temporal scanning. The output beams were combined and directed to a balanced photodetector (Femto, model HBPR-100M-60K-IN-FST) for detection of the interference signal. The resulting photocurrent was recorded using an oscilloscope (Keysight, model EXR404A, with 4 GHz bandwidth) for subsequent analysis. The delay line was moving periodically at rate of 5 Hz and an amplitude of 2 m and correspondingly produced a position signal that was sampled by the oscilloscope in synchronous with the signal from the balanced photodetector.
@@ 4.2 QD SOA fabrication
The active region of the gain medium we used comprised six layers of high-density, cm-2, InAs QDs grown by molecular beam epitaxy in the Stranski - Krastanow mode. The optical amplifier is fabricated as a 2 wide, 1.5 mm long ridge waveguide whose end facets were anti-reflection coated. The QDs exhibit a narrow photo-luminescence linewidth, measured at 10 K, of 17 meV for a single layer and 26 meV for the six-layer stack. This linewidth represents a record uniformity compared to all reported QDs in any material system. Details of the device fabrication and characterization are given in [banyoudeh2015high]. The emission power of the measured electro-luminescence spectra increases with applied bias but the spectral shape remains unchanged [khanonkin2017ultra].
@@ 4.3 Wigner function analysis
Histograms of the current at the output of the balanced photodetector, measured with a real time oscilloscope, is constructed from one thousand, periodic scans of the delay line with an amplitude of 2 m. Wigner functions are obtained from those histograms using the inverse Radon transform. This transform is intended for CW or narrow-spectrum signal. The spectral width of the 100 fs excitation pulses is roughly 3% of the optical frequency and hence, the assumption of narrow band signal is valid.
Acknowledgements The research was partially supported by the Israel Science Foundation under grant number 460/21 and by the Hellen Diller Quantum Center at Technion. IK acknowledges financial support from the Ministry of Innovation, Science and Technology of Israel.